# The inducibility of Turán graphs

### v1

@@ 1 Introduction
Results of Pippenger–Golumbic [27] (also Bollobás–Nara–Tachibana [6]) show that for bipartite Turán graphs $F$, the value of $I(F,n)$ is attained by the bipartite Turán graph $T_{2}(n)$. Using Zykov symmetrization [33] together with additional arguments, Brown–Sidorenko [7] showed that for every complete multipartite graph $F$, an extremal graph for $I(F,n)$ can always be chosen from the class of complete multipartite graphs; moreover, if $F$ is complete bipartite, then the number of parts can be taken to be at most two. Nevertheless, determining the number of parts and the ratios of part sizes in the extremal construction for a complete multipartite graph $F$ remains difficult in general. Consequently, the inducibility of complete multipartite graphs is still largely open.
@@ 1.1 Inducibility of almost balanced graphs
$\kappa_{F}\coloneqq\frac{\binom{\ell}{a_{1},\ldots,a_{r}}}{\mathrm{sym}(a_{1},\ldots,a_{r})},$ (2)
where $\mathrm{sym}(a_{1},\ldots,a_{r})$ denotes the number of distinct permutations of the multiset $\{\hskip-2.5pt\{\hskip 1.00006pta_{1},\ldots,a_{r}\hskip 1.00006pt\}\hskip-2.5pt\}$. Let $m_{r,\ell}$ denote the unique [^1] The uniqueness will be proved in Lemma 3.2. integer that maximizes the discrete function $f\colon[r,\infty)\to\mathbb{R}$ defined by
@@ 1.2 Inducibility of almost balanced graphs in $H$-free graphs
Given a graph $H$, we say that a graph $G$ is $H$-free if it does not contain $H$ as a (not necessarily induced) subgraph. Let $I_{H}(F,n)$ denote the maximum number of induced copies of $F$ in an $H$-free $n$-vertex graph, that is,
@@ 2 Preliminaries
Let $a_{1}\geq\cdots\geq a_{r}\geq 1$ be integers, and let $F=K_{a_{1},\ldots,a_{r}}$ be the complete $r$-partite graph with part sizes $a_{1},\ldots,a_{r}$. Suppose that $G$ is a complete multipartite graph whose non-singleton part sizes are $b_{1},\ldots,b_{k}$. Then
@@
In particular, if $\mathrm{sc}(F)=0$, then
@@
Therefore, by Lemmas 2.2 and 2.3, in determining $i(F)$ and $i_{k+1}(F)$, it suffices to restrict our attention to the class of complete multipartite graphs, which, in the limit, reduces to an optimization problem of homogeneous polynomials (see [7, Corollary 6]). Since we also aim to prove perfect stability, we need to consider a slightly more general polynomial (introduced in [22], see (6)) and prove the finiteness of the set of maximizers. For this, we require the following definitions, mostly taken from [22, Section 2].
@@
Note that for every $\bm{x}\in\mathrm{OPT}(F)$, the graph $G_{n,\bm{x}}$ is asymptotically extremal for the inducibility problem $I(F,n)$.
Similarly, it follows relatively straightforwardly from Lemma 2.3 and standard results from the theory of graphons (see [32, Section 1] for further details) that, for a complete multipartite graph $F=K_{a_{1},\ldots,a_{r}}$, we have
@@
Observe that $\mathbb{S}^{k-1}\subseteq\overline{\mathcal{P}}$ and the union $\bigcup_{k\geq 1}\mathbb{S}^{k-1}$ is the collection of sequences $\bm{x}\in\overline{\mathcal{P}}$ with $x_{0}=0$.
Define
@@
For convenience, we set $\mathbb{S}^{\infty}=\bigcup_{k\geq 1}\mathbb{S}^{k-1}$, $\mathrm{OPT}_{\infty}(F)=\mathrm{OPT}(F)$, and $i_{\infty}(F)=i(F)$.
For every $k\in\mathbb{N}_{+}$, define
@@ 3 Asymptotic results and uniqueness of the maximizer
We remark that the proof of the case $\ell\leq 2r-1$ in Theorem 3.1 is similar to that of [32, Theorem 1.6] (which deals with the case $\ell\leq 3r+1$), with one difference: for $k=\infty$, Yuster proved the uniqueness of the maximizer of (7), whereas we prove the uniqueness of the maximizer of (6) (which is slightly stronger but necessary for perfect stability). In contrast, our proof for the case $\ell\geq 2r$ differs substantially from Yuster’s. It relies on a quite nontrivial inequality (Proposition 3.7), whereas Yuster’s argument is based on properties of low-degree polynomials.
@@ 3.1 Basic properties of $f$ in (3) and the integer $m_{r,\ell}$
Let $I\subseteq\mathbb{R}$, and let $h\colon I\to\mathbb{R}$ be a function on $I$. We say that $h$ is unimodal if there exists a point $x_{0}\in I$ such that
@@
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\in[r+1,~2r-1]$ vertices (i.e. $\ell=a_{1}+\cdots+a_{r}$), and assume that $a_{1}\geq\cdots\geq a_{r}$. Let $m=m_{r,\ell}$ be the integer give by Lemma 3.2. By the Pigeonhole Principle, we have $a_{r}=1$. By the definition of almost balanced (see (1)), we have $\binom{a_{1}-1}{2}<1$, which means that $a_{1}\leq 2$. Thus we have
@@
Now, suppose $\ell=r+1$. Take any maximizer $\bm{x}\in\mathrm{OPT}_{k+1}(F)$. If $\bm{x}\neq\bm{k}$, then by Claim 3.6 (ii) we obtain another balanced maximizer $\bm{y}\in\mathrm{OPT}_{k+1}(F)$ with $|\mathrm{supp}(\bm{y})|<|\mathrm{supp}(\bm{x})|$. Consequently, $p_{F}(\bm{y})=p_{F}(\bm{t})$ for some $t<k$, contradicting $i_{k+1}(F)=p_{F}(\bm{k})>p_{F}(\bm{t})$. Therefore, we must have $\bm{x}=\bm{k}$, and again $\mathrm{OPT}_{k+1}(F)=\{\bm{k}\}$.
This completes the proof of Theorem 3.1 for $\ell\leq 2r-1$. ∎
@@
$\left(x^{k+1}+y^{k+1}-2\left(\tfrac{x+y}{2}\right)^{k+1}\right)\left((x+y)^{k}-x^{k}-y^{k}\right)\\ \leq\left(x^{k}+y^{k}-2\left(\tfrac{x+y}{2}\right)^{k}\right)\left((x+y)^{k+1}-x^{k+1}-y^{k+1}\right).$
@@
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\geq 2r$ vertices, that is, $\ell=a_{1}+\cdots+a_{r}$. Let $m=m_{r,\ell}$ be the integer give by Lemma 3.2. Assume that $a_{1}\geq\cdots\geq a_{r}$. Since $F$ is almost balanced, we have $a_{r}\geq 2$, otherwise by (1), we would have $a_{1}\leq 2$, which implies that $\ell\leq 2r-1$, a contradiction. Fix $k\in[r,\infty]$.
@@
It follows from the definition of $\mathbb{S}^{k-1}$ that $x_{0}=0$ if $k\neq\infty$. Thus we may assume that $k=\infty$. Suppose to the contrary that there exists $\bm{x}=(x_{1},x_{2},\ldots)\in\mathrm{OPT}(F)$ with $x_{0}>0$. Since in every realization $G_{n,\bm{{x}}}$ (Definition 2.4), there is no induced copy of $F$ containing vertices from $V_{0}$ (because $a_{1}\geq\cdots\geq a_{r}\geq 2$), we have
@@
It is clear that $|\mathrm{supp}(\bm{x})|\geq r$. Thus, $|\mathrm{supp}(\tilde{\bm{x}})|\geq r-2$, and hence, $B_{p,q}>0$ for all $\{p,q\}\subseteq[r]$.
Define
@@ 4.1 Preparations
In this subsection, we briefly review the sufficient conditions for perfect stability in the framework developed by [22]. Their framework applies to the broader class of “symmetrizable” functions, but here we focus specifically on the inducibility problem for complete multipartite graphs $F$. Very roughly, [22, Theorem 1.1] asserts that if the set of maximizers $\bm{x}$ satisfies certain strictness conditions, then the inducibility problem for $F$ is perfectly stable.
For convenience, we will use the following simplified variant of [22, Theorem 1.1], which is roughly [3, Theorem 8.2] (see also [4, Theorem 5.2]).
@@
Let $F$ be a complete multipartite graph with $\ell$ vertices. Suppose that $\mathrm{OPT}(F)$ contains a unique sequence $\bm{x}$ with $x_{0}=0$ (i.e. $\sum_{i\in\mathbb{N}_{+}}x_{i}=1$) and $t\coloneqq|\mathrm{supp}(\bm{x})|<\infty$. Suppose that there exist constants $\varepsilon>0$ and $N_{0}$ such that the following holds for the complete multipartite graph $G\coloneqq G_{n,\bm{x}}$ with parts $V_{1},\ldots,V_{t}$ for every $n\geq N_{0}\colon$
@@
    $I(F,G)-I(F,G\oplus xy)\geq\varepsilon n^{\ell-2},$
    where $G\oplus xy$ denotes the graph with vertex set $V(G)$ and edge set $E(G)\bigtriangleup\{xy\}$.
@@
    For every $A\subseteq[t]$ with $|A|\neq t-1$ and every $A^{*}\subseteq[t]$ with $|A^{*}|=t-1$,
    $I(F,G_{A^{*}})-I(F,G_{A})\geq\varepsilon n^{\ell-1},$
    where for a subset $B\subseteq[t]$, $G_{B}$ denotes the graph obtained from $G$ by adding a new vertex adjacent precisely to $\bigcup_{i\in B}V_{i}$.
@@
Next, we derive refined estimates for $m_{r,\ell}$, the integer given by Lemma 3.2.
@@
Let $\ell>r\geq 2$ be integers and let $m=m_{r,\ell}$ be the unique integer from Lemma 3.2. Then the following inequalities hold:
@@
    $m>\max\left\{\frac{\ell(r-1)}{2(\ell-r)},~r-1\right\}$.
@@
    $m<\min\left\{\frac{\ell(\ell-1)}{2(\ell-r)},~\frac{r}{\alpha}\right\}$, where $\alpha\coloneqq\alpha(\ell/r)\in(1-r^{2}/\ell^{2},1)$ is the unique positive root of the equation $\mathrm{e}^{(\ell/r)\cdot x}(1-x)=1$. In particular, $m<\frac{r\ell^{2}}{\ell^{2}-r^{2}}$.
@@
First, we prove (i). Since we trivially have $m\geq r$, it suffices to show that $m>\frac{\ell(r-1)}{2(\ell-r)}$. Note that $\frac{\ell(r-1)}{2(\ell-r)}$ is decreasing in $\ell>r$, and when $\ell=2r$, we have $\frac{\ell(r-1)}{2(\ell-r)}<r$. Thus we may assume that $r<\ell<2r$.
Since $f(x)$ (as defined in (3)) is unimodal, the optimality of $m$ implies that $f(m)>f(m+1)$. Simplifying this inequality yields
@@
$\left(1-\frac{2(\ell-r)}{r\ell+\ell-2r}\right)^{\ell}\geq 1-\frac{2r(\ell-r)}{r\ell+\ell-2r}.$
Taking logarithms, it simplifies to showing that
$\ell\ln\ell+(\ell-1)\ln(r-1)-(\ell-1)\ln((r+1)\ell-2r)-\ln(2r-\ell)\geq 0.$ (23)
@@
Straightforward calculations show that $\vartheta(r)=\vartheta^{\prime}(r)=0$ and
$\vartheta^{\prime\prime}(x)=\frac{4r^{2}(x-r)\bigl((r+2)x-4r\bigr)}{\bigl((r+1)x-2r\bigr)^{2}(2r-x)^{2}x}.$
@@
Since the second derivative $g^{\prime\prime}(x)=\ell(\ell-1)(1-x)^{\ell-2}\geq 0$, the function $g(x)$ is convex. Note that $g(0)=0$ and $g\left(\frac{1}{m+1}\right)<0$. Combining this with Claim 4.3, we obtain $\frac{1}{1+m}<\frac{2(\ell-r)}{r\ell+\ell-2r}$, which implies that $m>\frac{\ell(r-1)}{2(\ell-r)}$. This proves (i).
It remains to prove (ii). Let $t=\ell/r$. First we show that the real number $\alpha$ defined in the lemma lies in the interval $(1-1/t^{2},1)$. Define $h_{t}(x)\coloneqq\mathrm{e}^{tx}(1-x)-1$. Note that the derivative $h_{t}^{\prime}(x)=\mathrm{e}^{tx}(t-1-tx)$ is positive when x<11/tx<1-1/t and negative when x>11/tx>1-1/t. Thus, $h_{t}(x)$ is unimodal on $(0,\infty)$. Since $h_{t}(0)=0$ and $h_{t}(1)=-1$, the equation $h_{t}(x)=1$ has a unique positive root, which lies in the interval $(0,1)$.
@@
$h_{t}(1-1/t^{2})=\frac{\mathrm{e}^{t-\frac{1}{t}}}{t^{2}}-1.$
Let $u(t)\coloneqq\mathrm{e}^{t-\frac{1}{t}}/t^{2}$ for $t>1$. Since the derivative $u^{\prime}(t)=u(t)(1-1/t)^{2}$ is positive, the function $u(t)$ is strictly increasing. Consequently, $h_{t}(1-1/t^{2})=u(t)-1>u(1)-1=0$, as desired.
@@
Since $f(x)$ (as defined in (3)) is unimodal, the optimality of $m$ implies that f(m)>f(m1)f(m)>f(m-1). Simplifying this inequality yields
@@
In other words, we have $h_{t}(\theta)<0$. Since $h_{t}(x)$ is unimodal and $h_{t}(0)=0$, we have $\theta>\alpha$, which implies that $m<r/\alpha$. This proves (ii), completing the proof of Lemma 4.2. ∎

### v2

@@ 1 Introduction
Results of Pippenger–Golumbic [27] (also Bollobás–Nara–Tachibana [6]) show that for bipartite Turán graphs $F$, the value of $I(F,n)$ is attained by bipartite Turán graphs $T_{2}(n)$. Using Zykov symmetrization [33] together with additional arguments, Brown–Sidorenko [7] showed that for every complete multipartite graph $F$, an extremal graph for $I(F,n)$ can always be chosen from the class of complete multipartite graphs; moreover, if $F$ is complete bipartite, then the number of parts can be taken to be at most two. Nevertheless, determining the number of parts and the ratios of part sizes in the extremal construction for a complete multipartite graph $F$ remains difficult in general. Consequently, the inducibility of complete multipartite graphs is still largely open.
@@ 1.1 Inducibility of almost balanced graphs
$\kappa_{F}\coloneqq\frac{\binom{\ell}{a_{1},\ldots,a_{r}}}{\mathrm{sym}(a_{1},\ldots,a_{r})}=\frac{\ell!}{a_{1}!\cdots a_{r}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}.$ (2)
Here, $\mathrm{sym}(a_{1},\ldots,a_{r})$ denotes the size of the symmetry group of the multiset $\{\hskip-2.5pt\{\hskip 1.00006pta_{1},\ldots,a_{r}\hskip 1.00006pt\}\hskip-2.5pt\}$. In other words, if we let $b_{1},\ldots,b_{t}$ denote the distinct elements of the multiple set $\{\hskip-2.5pt\{\hskip 1.00006pta_{1},\ldots,a_{r}\hskip 1.00006pt\}\hskip-2.5pt\}$, occurring with multiplicities $r_{1},\ldots,r_{t}$, respectively, then
$\mathrm{sym}(a_{1},\ldots,a_{r})=r_{1}!\cdots r_{t}!.$
Let $m_{r,\ell}$ denote the unique [^1] The uniqueness will be proved in Lemma 3.2. integer that maximizes the discrete function $f\colon[r,\infty)\to\mathbb{R}$ defined by
@@ 1.2 Inducibility of almost balanced graphs in $H$-free graphs
Given a graph $H$, we say that a graph $G$ is $H$-free if it does not contain $H$ as a (not necessarily induced) subgraph. Let $I_{H}(F,n)$ denote the maximum number of induced copies of $F$ in an $H$-free $n$-vertex graph, i.e.,
@@ 2 Preliminaries
Let $k$ be a positive integer. For every real number $z$, define $\binom{z}{k}=(z)_{k}/k!$, where
$(z)_{k}\coloneqq\begin{cases}0,&\quad\text{if}\quad z\leq k-1,\\ z\cdots(z-k+1),&\quad\text{if}\quad z>k-1.\end{cases}$
@@
Let $F=K_{a_{1},\ldots,a_{r}}$ be the complete $r$-partite graph with part sizes $a_{1}\geq\cdots\geq a_{r}\geq 1$. Suppose that $G$ is a complete multipartite graph whose non-singleton part sizes are $b_{1},\ldots,b_{k}$. Then
@@
In particular, if $\mathrm{sc}(F)=0$ or $\mathrm{sc}(G)=0$, then
@@
A more convenient version of Fact 2.1 for the case in which each part of $G$ has approximately the same size and each part is large, which will be useful in the proof of perfect stability, is as follows.
Fact 2.2.
Let $F=K_{a_{1},\ldots,a_{r}}$ be the complete $r$-partite graph with part sizes $a_{1}\geq\cdots\geq a_{r}\geq 1$. Suppose that $G$ is a complete $m$-partite graph in which each part has size $(1+o(1))n$, where $n$ is large. Then
$I(F,G)=(1+o(1))\frac{(m)_{r}}{a_{1}!\cdots a_{r}!\cdot\mathrm{sym}(a_{1},\ldots,a_{r})}n^{\ell}.$ (5)
@@
Therefore, by Lemmas 2.3 and 2.4, in determining $i(F)$ and $i_{k+1}(F)$, it suffices to restrict our attention to the class of complete multipartite graphs, which, in the limit, reduces to an optimization problem of homogeneous polynomials (see [7, Corollary 6]). Since we also aim to prove perfect stability, we need to consider a slightly more general polynomial (introduced in [22], see (7)) and prove the finiteness of the set of maximizers. For this, we require the following definitions, mostly taken from [22, Section 2].
@@
So for every $\bm{x}\in\mathrm{OPT}(F)$, the graph $G_{n,\bm{x}}$ is asymptotically extremal for the inducibility problem $I(F,n)$.
It follows relatively straightforwardly from Lemma 2.4 and standard results from the theory of graphons (see [32, Section 1] for further details) that, for a complete multipartite graph $F=K_{a_{1},\ldots,a_{r}}$, we have
@@
Observe that $\mathbb{S}^{k-1}\subseteq\overline{\mathcal{P}}$ and the union $\bigcup_{k\geq 1}\mathbb{S}^{k-1}$ is the collection of sequences $\bm{x}\in\overline{\mathcal{P}}$ with $x_{0}=0$. Define
@@
For convenience, set $\mathbb{S}^{\infty}=\bigcup_{k\geq 1}\mathbb{S}^{k-1}$, $\mathrm{OPT}_{\infty}(F)=\mathrm{OPT}(F)$, and $i_{\infty}(F)=i(F)$. For every $k\in\mathbb{N}_{+}$, define
@@
The following simple property of almost balanced complete multipartite graphs will be used multiple times.
Fact 2.7.
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph with $a_{1}\geq\cdots\geq a_{r}$, and let $\ell\coloneqq a_{1}+\cdots+a_{r}$.
  1. (i)
    If $\ell\in[r+1,2r-1]$, then $a_{1}=\cdots=a_{\ell-r}=2$ and $a_{\ell-r+1}=\cdots=a_{r}=1$.
  2. (ii)
    If $\ell\geq 2r$, then $a_{r}\geq 2$.
Proof.
Suppose that $\ell\in[r+1,2r-1]$. By the Pigeonhole Principle, we have $a_{r}=1$. By the definition of almost balanced (see (1)), we have $\binom{a_{1}-1}{2}<1$, which implies that $a_{1}\leq 2$. Therefore,
Now suppose that $\ell\geq 2r$. If $a_{r}=1$, then the definition of almost balanced implies that $\binom{a_{1}-1}{2}<1$, which forces $a_{1}\leq 2$. Consequently, $\ell\leq 2r-1$, which is a contradiction. ∎
@@ 3 Asymptotic results and uniqueness of the maximizer
We remark that the proof of the case $\ell\leq 2r-1$ in Theorem 3.1 is similar to that of [32, Theorem 1.6] (which works for $\ell\leq 3r+1$), with one difference: for $k=\infty$, Yuster proved the uniqueness of the maximizer of (8), whereas we prove the uniqueness of the maximizer of (7) (slightly stronger but necessary for perfect stability). In contrast, our proof for the case $\ell\geq 2r$ differs substantially from Yuster’s. It relies on a quite nontrivial inequality (Proposition 3.7), whereas Yuster’s argument is based on properties of low-degree polynomials.
@@ 3.1 Basic properties of $f$ in (3) and the integer $m_{r,\ell}$
Let $h\colon I\to\mathbb{R}$ be a function on $I\subseteq\mathbb{R}$. We say that $h$ is unimodal if there exists a point $x_{0}\in I$ such that
@@
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\in[r+1,~2r-1]$ vertices (i.e. $\ell=a_{1}+\cdots+a_{r}$), and assume that $a_{1}\geq\cdots\geq a_{r}$. Let $m=m_{r,\ell}$ be the integer give by Lemma 3.2. By Fact 2.7 (i), we have $a_{1}=\cdots=a_{\ell-r}=2$ and $a_{\ell-r+1}=\cdots=a_{r}=1$.
@@
Now, suppose $\ell=r+1$. Take an arbitrary maximizer $\bm{x}\in\mathrm{OPT}_{k+1}(F)$. If $\bm{x}\neq\bm{k}$, then by Claim 3.6 (ii) we obtain another balanced maximizer $\bm{y}\in\mathrm{OPT}_{k+1}(F)$ with $|\mathrm{supp}(\bm{y})|<|\mathrm{supp}(\bm{x})|$. Consequently, $p_{F}(\bm{y})=p_{F}(\bm{t})$ for some $t<k$, contradicting $i_{k+1}(F)=p_{F}(\bm{k})>p_{F}(\bm{t})$. Therefore, we must have $\bm{x}=\bm{k}$, and again $\mathrm{OPT}_{k+1}(F)=\{\bm{k}\}$. This completes the proof of Theorem 3.1 for $\ell\leq 2r-1$. ∎
@@
$\left(x^{k+1}+y^{k+1}-2\left(\tfrac{x+y}{2}\right)^{k+1}\right)\left((x+y)^{k}-x^{k}-y^{k}\right)\leq\left(x^{k}+y^{k}-2\left(\tfrac{x+y}{2}\right)^{k}\right)\left((x+y)^{k+1}-x^{k+1}-y^{k+1}\right).$
@@
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell\geq 2r$ vertices, that is, $\ell=a_{1}+\cdots+a_{r}$. Let $m=m_{r,\ell}$ be the integer give by Lemma 3.2. Assume that $a_{1}\geq\cdots\geq a_{r}$. By Fact 2.7 (ii), we have $a_{r}\geq 2$. Fix an integer $k\in[r,\infty]$.
@@
It follows from the definition of $\mathbb{S}^{k-1}$ that $x_{0}=0$ if $k\neq\infty$. Thus we may assume that $k=\infty$. Suppose to the contrary that there exists $\bm{x}=(x_{1},x_{2},\ldots)\in\mathrm{OPT}(F)$ with $x_{0}>0$. Since in every realization $G_{n,\bm{{x}}}$ (see Definition 2.5), no induced copy of $F$ can contain vertices from $V_{0}$ (because $a_{1}\geq\cdots\geq a_{r}\geq 2$), we have
@@
It is clear that $|\mathrm{supp}(\bm{x})|\geq r$. Thus, $|\mathrm{supp}(\tilde{\bm{x}})|\geq r-2$, and hence, $B_{p,q}>0$ for all $\{p,q\}\subseteq[r]$. Define
@@ 4.1 Preparations
In this subsection, we briefly review the sufficient conditions for perfect stability in the framework developed by [22]. Their framework applies to the broader class of “symmetrizable” functions, but here we focus specifically on the inducibility problem for complete multipartite graphs $F$. Very roughly, [22, Theorem 1.1] asserts that if $\mathrm{OPT}(F)$ is finite and each maximizer $\bm{x}\in\mathrm{OPT}(F)$ satisfies certain strictness conditions, then the inducibility problem for $F$ is perfectly stable.
For convenience, we will use the following simplified variant of [22, Theorem 1.1] (see also [3, Theorem 8.2]).
For a graph $G$ and a pair $\{x,y\}\subseteq V(G)$ of vertices, let $G\oplus xy$ denote the graph with vertex set $V(G)$ and edge set $E(G)\bigtriangleup\{xy\}$. That is, $G\oplus\{xy\}$ is obtained by adding the edge $\{x,y\}$ if it is not present in $G$, and by removing it otherwise. Suppose that $G$ is a complete $t$-partite graph with parts $V_{1},\ldots,V_{t}$. For every $A\subseteq[t]$, let $G_{A}$ denote the graph obtained from $G$ by adding a new vertex that is adjacent precisely to all vertices in $\bigcup_{i\in A}V_{i}$.
@@
Let $F$ be a complete multipartite graph with $\ell$ vertices. Suppose that $\mathrm{OPT}(F)$ contains a unique sequence $\bm{x}$ with $x_{0}=0$ (i.e. $\sum_{i\in\mathbb{N}_{+}}x_{i}=1$) and $t\coloneqq|\mathrm{supp}(\bm{x})|<\infty$. Suppose that there exist constants $\varepsilon>0$ and $N_{0}$ such that for every $n\geq N_{0}$ the following statements hold for the $n$-vertex realization $G\coloneqq G_{n,\bm{x}}$ with parts $V_{1},\ldots,V_{t}\colon$
@@
    $I(F,G)-I(F,G\oplus xy)\geq\varepsilon n^{\ell-2}.$
@@
    For every $A\subseteq[t]$ with $|A|\neq t-1$, we have
    $\min\big\{I(F,G_{A^{\ast}})\colon A^{\ast}\subseteq[t]\text{ and }|A^{\ast}|=t-1\big\}-I(F,G_{A})\geq\varepsilon n^{\ell-1}.$
@@
We also need the following refined estimates for $m_{r,\ell}$, the integer given by Lemma 3.2.
@@
Let $\ell>r\geq 2$ be integers and let $m=m_{r,\ell}$. Then
@@
    $m>\max\left\{\frac{\ell(r-1)}{2(\ell-r)},~r-1\right\}$, and
@@
    $m<\min\left\{\frac{\ell(\ell-1)}{2(\ell-r)},~\frac{r}{\alpha}\right\}$, where $\alpha\coloneqq\alpha(\ell/r)\in(1-r^{2}/\ell^{2},1)$ is the unique positive real root of the equation $\mathrm{e}^{(\ell/r)\cdot x}(1-x)=1$. In particular, $m<\frac{r\ell^{2}}{\ell^{2}-r^{2}}$.
@@
First, we prove (i). Since $m\geq r$ holds trivially, it suffices to show that $m>\frac{\ell(r-1)}{2(\ell-r)}$, equivalently, $\frac{1}{m+1}<\frac{2(\ell-r)}{r\ell+\ell-2r}$. Note that $\frac{\ell(r-1)}{2(\ell-r)}$ is decreasing in $\ell>r$, and when $\ell=2r$, we have $\frac{\ell(r-1)}{2(\ell-r)}<r$. Thus we may assume that $r<\ell<2r$.
By Lemma 3.2, the discrete function $f$ defined in (3) is unimodal. Therefore, the optimality of $m$ implies that $f(m)>f(m+1)$, which yields
@@
Straightforward calculations show that the second derivative of $g$ is $g^{\prime\prime}(x)=\ell(\ell-1)(1-x)^{\ell-2}$, which is nonnegative for $x\in[0,1]$. Thus, the function $g$ is convex on the interval $[0,1]$. Since $g(0)=0$ and $g\left(\frac{1}{m+1}\right)<0$, to show that $\frac{1}{m+1}<\frac{2(\ell-r)}{r\ell+\ell-2r}$, it suffices to prove the following claim.
@@
$\left(\frac{(r-1)\ell}{(r+1)\ell-2r}\right)^{\ell}=\left(1-\frac{2(\ell-r)}{r\ell+\ell-2r}\right)^{\ell}\geq 1-\frac{2r(\ell-r)}{r\ell+\ell-2r}=\frac{(2r-\ell)(r-1)}{(r+1)\ell-2r}.$
Taking logarithms on both sides and then rearranging the terms, this simplifies to showing that
$\ell\ln\ell+(\ell-1)\ln(r-1)-(\ell-1)\ln\big((r+1)\ell-2r\big)-\ln(2r-\ell)\geq 0.$ (24)
@@
Straightforward calculations show that
$\vartheta(r)=\vartheta^{\prime}(r)=0\quad\text{and}\quad\vartheta^{\prime\prime}(x)=\frac{4r^{2}(x-r)\bigl((r+2)x-4r\bigr)}{\bigl((r+1)x-2r\bigr)^{2}(2r-x)^{2}x}.$
@@
It remains to prove (ii). Let $t\coloneqq\ell/r$ and $h_{t}(x)\coloneqq\mathrm{e}^{tx}(1-x)-1$. First we show that the real number $\alpha$ defined in the lemma lies in the interval $(1-1/t^{2},1)$. Note that the derivative $h_{t}^{\prime}(x)=\mathrm{e}^{tx}(t-1-tx)$ is positive when x<11/tx<1-1/t and negative when x>11/tx>1-1/t. Thus, $h_{t}$ is unimodal on $(0,\infty)$. Since $h_{t}(0)=0$ and $h_{t}(1)=-1$, the equation $h_{t}(x)=1$ has a unique positive root (i.e. $\alpha$), which lies in the interval $(0,1)$.
@@
$h_{t}(1-1/t^{2})=\frac{1}{t^{2}}\mathrm{e}^{t-\frac{1}{t}}-1.$
Let $u(t)\coloneqq\mathrm{e}^{t-\frac{1}{t}}/t^{2}$. Since the derivative $u^{\prime}(t)=u(t)(1-1/t)^{2}$ is positive for $t>1$, the function $u$ is strictly increasing on $(1,\infty)$. Consequently, $h_{t}(1-1/t^{2})=u(t)-1>u(1)-1=0$ for $t>1$, as desired. This proves that $\alpha\in(1-1/t^{2},1)$.
@@
By Lemma 3.2, the discrete function $f$ defined in (3) is unimodal. Therefore, the optimality of $m$ implies that f(m)>f(m1)f(m)>f(m-1), which yields
@@
In other words, we have $h_{t}(\theta)<0$. Since $h_{t}$ is unimodal on $(0,\infty)$ and $h_{t}(0)=0$, we have $\theta>\alpha$, which implies that $m<r/\alpha$. This proves (ii), and thus completing the proof of Lemma 4.2. ∎